OpenStudy (anonymous):

What value of c makes the polynomial below a perfect square ? x^2+14x+c

zepdrix (zepdrix):

x^2 +bx + c We have an equation of this form :D divide your b term by 2. Then square the result, and that is c! Example:\[x^2 +16x + c\]\[\frac{ 16 }{ 2 }=8\]\[8^2=64\]\[x^2+16x + 64 = (x+8)^2\] A perfect square! :)

OpenStudy (anonymous):

So its 2 ?

OpenStudy (anonymous):

Um I actually said X^2 +14x+c ... 14 not 16 . but anyways . wouldn't 49 be the correct answer ? i'm having trouble checking myself.

zepdrix (zepdrix):

yes, i was just giving an example :) 49! gj